# Multimodal Embodied Interactive Agent for Cafe Scene

### v1

@@ Abstract
With the surge in the development of large language models, embodied intelligence has attracted increasing attention. Nevertheless, prior works on embodied intelligence typically encode scene or historical memory in an unimodal manner, either visual or linguistic, which complicates the alignment of the model’s action planning with embodied control. To overcome this limitation, we introduce the Multimodal Embodied Interactive Agent (MEIA), capable of translating high-level tasks expressed in natural language into a sequence of executable actions. Specifically, we propose a novel Multimodal Environment Memory (MEM) module, facilitating the integration of embodied control with large models through the visual-language memory of scenes. This capability enables MEIA to generate executable action plans based on diverse requirements and the robot’s capabilities. We conduct experiments in a dynamic virtual cafe environment, utilizing multiple large models through zero-shot learning, and carefully design scenarios for various situations. The experimental results showcase the promising performance of our MEIA in various embodied interactive tasks.
@@ 1 Introduction
Picture yourself entering a coffee shop staffed by robots. Upon recognizing your arrival and that of your friends, a robot approaches, greets you, and guides you to suitable seats based on the group size. Subsequently, it proceeds to the bar, takes your coffee and dessert orders, and then prepares your order. Meanwhile, you request the robot to adjust the environment by lowering the curtains and decreasing the air conditioning due to the rising outside temperature and warmth inside the cafe. This scenario illustrates the agent’s responsibilities in vision-and-language navigation, instruction decomposition, and planning, as depicted in Fig.1. These tasks demand diverse capabilities, including visual recognition Liu et al. (2018c, 2023d); Yan et al. (2023), target navigation Liu et al. (2022b), language comprehension Liu et al. (2023c); Wei et al. (2023), commonsense reasoning Liu et al. (2023b); Chen et al. (2023), task planning Zhu et al. (2022); Wang et al. (2023a); Lin et al. (2023), and object manipulation Tang et al. (2023). Termed embodied AI, it stands as a prominent research focus in the realms of artificial intelligence and robotics. Unlike conventional AI learning from datasets comprised mainly of internet-curated images Liu et al. (2016, 2019); Xu et al. (2022), videos Liu et al. (2022a, 2021, 2018a, 2018b), or text Liu et al. (2023e), embodied AI involves learning from self-centered data, akin to human experiences Duan et al. (2022). Agents must collect vital visual information and engage with the physical world based on real-time observations in embodied AI tasks.
@@
To address the above challenges, we propose a novel embodied model MEIA based on LLMs and VLMs, which excels in performing a variety of embodied AI tasks such as embodied question answering and embodied control in a physical environment. Our MEIA comprises three essential modules that work collaboratively, including i) vision module utilized to generate environmental language memory and environmental image memory based on the surroundings, ii) control module used to dialogue with guests and perform embodied operations such as seating guidance and cleaning up, iii) large model responsible for reasoning and action planning. Since a good environmental memory storage method aids agents in remembering the environment and making correct decisions to facilitate task completion, we propose a novel Multi-modal Environmental Memory (MEM) module incorporating object descriptions and environmental pictures. To enhance the accuracy and feasibility of model output, we also meticulously designed prompts to guide the LLMs and VLMs. We evaluate MEIA in the simulator provided by Dataa Robotics, which provides a cafe scene that is close to reality and achieves abundant human-machine interaction design. The contributions can be summarized as follows:
@@
  • Our MEIA exhibits robust zero-shot learning capabilities and is capable of completing the embodied task with nearly perfect performance.
@@ 3.2 Robot Configuration
In addition, the robot can perform complex interactive operations such as turning on/off lights, moving tables and chairs, pushing carts, using vacuum cleaners, using coffee machines, and operating air conditioning control panels, as shown in Fig.1.
@@ 4.1 Multimodal Environment Memory
The embodied control and the large model cannot be directly interacted due to the domain gap. The integration of embodied control and the large model has always been the focus of embodied AI. Previous works Song et al. (2023); Zhou et al. (2023) usually implement memory of scenes or historical trajectories in a single-modal form of visual or natural language modality, making the plans given by the large model difficult to execute by embodied agents. Recently, some works attempt to delineate the scope of action plans generated by large models to achieve the executability of the plans Wake et al. (2023a). To address this issue, we propose the multimodal environment memory (MEM) module (Fig.2), which stores the environment image, the names and coordinates of key items in the scene recorded in natural language as memory, serving as a guidance for the large model to achieve highly executable action plans under diverse needs. The details of the construction of MEM are as follows.
The visual interface of the robot can obtain different modalities of images, including RGB images, depth images, and segmentation images. RGB images are directly stored as the image memory. Segmentation images set the pixels corresponding to different key items in the scene to their ID values to achieve item segmentation. We use the K-Means algorithm to calculate the pixel coordinates of the key items. The algorithm is shown in Algorithm 1.
@@
Obtaining the ginger coordinate system coordinates $(x_{g},y_{g},z_{g})$, the ginger coordinate system is opposite to the x-axis of the world coordinate system, and has Euler angles $\theta_{g}$ and translation vector $T_{g}$.
@@
By combining object IDs and coordinates with environment images, we obtain multimodal environment memory. The MEM is designed as the global memory so that the robot can update multimodal memory when it calls the visual interface, to achieve flexible response for complex environment.
@@ 4.2 Embodied Control
Our design for embodied control capability is versatile, allowing the robot to possess a range of abilities. It is not restricted to a single type of skill but is equipped to successfully accomplish a variety of tasks in complex scenarios. Our embodied control module not only provides “tools” for the large model module but also integrates multimodal environmental memory to execute the action sequence planned by the large model to complete the task.
@@ 4.3 Large Model
VLM. GPT-4 Turbo with vision has superior capability in both visual and language understanding compared to open-source models. In our experiments, it can form spatial understanding of complex images and offer independent judgments, thereby generating content of high quality. Therefore, we use GPT-4 Turbo with vision to implement guest seating guidance task (visual language navigation), spatial-aware dialogue interaction task (understanding the space inside the cafe and completing dialogue requirements), and control guidance task (determining which action to execute).
@@ 5 Experimental Results
We design a plot pipeline for our solution based on the reality, systematically evaluating and comparing various models for different tasks. The details are outlined below.
@@ 6 Conclusion
In this work, we propose the MEIA for accomplishing embodied tasks. In the physical environment, the agent gathers egocentric environmental information, executes embodied control and completes embodied tasks. We innovatively introduce multi-modal memory to store global memory and update it in real time, thereby integrating embodied control with large models. Experiments indicate that MEIA demonstrates the promising ability to solve various embodied tasks.
Future work. i) Enhance scenario diversity. While we have the flexibility to adapt the cafe environment as needed, we intend to introduce additional scenarios for a more comprehensive evaluation. ii) Improve embodied control. In subsequent research, we aim to further develop the robot’s ability to precisely and diversely manipulate its surroundings.

### v2

# MEIA: Towards Realistic Multimodal Interaction and Manipulation for Embodied Robots

@@ Abstract
With the surge in the development of large language models, embodied intelligence has attracted increasing attention. Nevertheless, prior works on embodied intelligence typically encode scene or historical memory in an unimodal manner, either visual or linguistic, which complicates the alignment of the model’s action planning with embodied control. To overcome this limitation, we introduce the Multimodal Embodied Interactive Agent (MEIA), capable of translating high-level tasks expressed in natural language into a sequence of executable actions. Specifically, we propose a novel Multimodal Environment Memory (MEM) module, facilitating the integration of embodied control with large models through the visual-language memory of scenes. This capability enables MEIA to generate executable action plans based on diverse requirements and the robot’s capabilities. Furthermore, we construct an embodied question answering dataset based on a dynamic virtual cafe environment with the help of the large language model. In this virtual environment, we conduct several experiments, utilizing multiple large models through zero-shot learning, and carefully design scenarios for various situations. The experimental results showcase the promising performance of our MEIA in various embodied interactive tasks.
@@ 1 Introduction
Picture yourself entering a coffee shop staffed by robots. Upon recognizing your arrival and that of your friends, a robot approaches, greets you, and guides you to suitable seats based on the group size. Subsequently, it proceeds to the bar, takes your coffee and dessert orders, and then prepares your order. Meanwhile, you request the robot to adjust the environment by lowering the curtains and decreasing the air conditioning due to the rising outside temperature and warmth inside the cafe. Additionally, it is able to respond to questions you asked based on the environment. This scenario illustrates the agent’s responsibilities in vision-and-language navigation, instruction decomposition and planning, as depicted in Fig.1. These tasks demand diverse capabilities, including visual recognition Liu et al. (2018c, 2023d); Yan et al. (2023), target navigation Liu et al. (2022b), language comprehension Liu et al. (2023c); Wei et al. (2023), commonsense reasoning Liu et al. (2023b); Chen et al. (2023), task planning Zhu et al. (2022); Wang et al. (2023a); Lin et al. (2023), and object manipulation Tang et al. (2023). Termed embodied AI, it stands as a prominent research focus in the realms of artificial intelligence and robotics. Unlike conventional AI learning from datasets comprised mainly of internet-curated images Liu et al. (2016, 2019); Xu et al. (2022), videos Liu et al. (2022a, 2021, 2018a, 2018b), or text Liu et al. (2023e), embodied AI involves learning from self-centered data, akin to human experiences Duan et al. (2022). Agents must collect vital visual information and engage with the physical world based on real-time observations in embodied AI tasks.
@@
To address the above challenges, we propose a novel embodied model MEIA based on LLMs and VLMs, which excels in performing a variety of embodied AI tasks such as embodied question answering and embodied control in a physical environment. Our MEIA comprises three essential modules that work collaboratively, including i) vision module utilized to generate environmental language memory and environmental image memory based on the surroundings, ii) control module used to dialogue with guests and perform embodied operations such as seating guidance and cleaning up, iii) large model responsible for reasoning and action planning. Since a good environmental memory storage method aids agents in remembering the environment and making correct decisions to facilitate task completion, we propose a novel Multi-modal Environmental Memory (MEM) module incorporating object descriptions and floor plan of the environment. To further enhance the accuracy and feasibility of model output, we meticulously designed prompts to guide the LLMs and VLMs. In addition, we also build an embodied question answering dataset with a thousand pieces of data based on the simulator provided by Dataa Robotics, which provides a cafe scene that is close to reality and achieves abundant human-machine interaction design. We evaluate our MEIA in the simulator and on the dataset we constructed. The contributions can be summarized as follows:
@@
  • We utilize a large language model to effectively generate data with high quality, constructing an embodied question answering dataset.
  • Our MEIA exhibits robust zero-shot learning capabilities and is capable of completing the embodied task with good performance.
@@ 3.2 Robot Configuration
In addition, the robot can display the content of the conversation. It can also perform complex interactive operations such as turning on/off lights, moving tables and chairs, pushing carts, using vacuum cleaners, using coffee machines, and operating air conditioning control panels, as shown in Fig.1.
@@ 4.1 Multimodal Environment Memory
The embodied control and the large model cannot be directly interacted due to the domain gap. The integration of embodied control and the large model has always been the focus of embodied AI. Previous works Song et al. (2023); Zhou et al. (2023) usually implement memory of scenes or historical trajectories in a single-modal form of visual or natural language modality, making the plans given by the large model difficult to execute by embodied agents. Recently, some works attempt to delineate the scope of action plans generated by large models to achieve the executability of the plans Wake et al. (2023a). To address this issue, we propose the multimodal environment memory (MEM) module (Fig.2), which stores the names and coordinates of key items in the scene recorded in natural language and the two-dimensional plan of the environment as memory, serving as a guidance for the large model to achieve highly executable action plans under diverse needs. The details of the construction of MEM are as follows.
Environmental Language Memory. The visual interface of the robot can obtain different modalities of images, including RGB images, depth images, and segmentation images. Segmentation images set the pixels corresponding to different key items in the scene to their ID values to achieve item segmentation. We use the K-Means algorithm to calculate the pixel coordinates of the key items. The algorithm is shown in Algorithm 1.
@@
When obtaining the coordinates $(x_{g},y_{g},z_{g})$ of the ginger coordinate system, it should be noted that the ginger coordinate system is opposite to the x-axis of the world coordinate system. Additionally, it has Euler angles $\theta_{g}$ and a translation vector $T_{g}$.
@@
Environmental Image Memory. The environmental image memory is an environmental plane map obtained by further projecting the 3D point cloud image constructed by the robot using the visual information gathered during the exploration process.
The robot obtains environmental visual information by calling the RGBD camera to capture RGB images and depth images. For the RGB image and depth image corresponding to the same viewing angle, with the help of Open3D, an open-source library that supports rapid processing of 3D data, a three-dimensional point cloud with color information is constructed based on the internal parameter information of the camera. However, directly superimposing 3D point clouds obtained from multiple sets of RGBD images will cause all point cloud information to overlap and block each other. Therefore, fusing the 3D point clouds obtained from each set of images also requires calculations similar to the world coordinates of the objects, applying rotation and translation operations based on the camera’s posture.
Additionally, outliers in the point cloud are identified and removed based on statistical analysis. For each point in the point cloud, we select the closest $n$ points and calculate their average distance from the point. Simultaneously, we calculate the global average distance $dis_{g}$ and global standard deviation $std_{g}$, which represent the average distance from each point in the point cloud to any other point and the standard deviation of these distances, respectively. If a point’s average distance is greater than $dis_{g}+std_{r}*std_{g}$, it is considered an outlier, where $std_{r}$ specifies the multiple of the standard deviation.
Although the 3D point cloud contains richer and more accurate information than 2D images, processing and analyzing three-dimensional images directly can be challenging for large models. As a result, we project the 3D point cloud onto the plane where the floor is located. The environmental plane map can to some extent display the exploration status of the current environment, providing an overall understanding of the cafe’s environment and determining whether the environment has been fully explored. As the exploration progresses, the 3D point cloud is enriched and the environmental plane map contains more information, as shown in Fig.3.
Figure 3: Construction process of an environmental floor plan.
By combining object IDs and coordinates with environmental plane map, we obtain multimodal environment memory. The MEM is designed as the global memory so that the robot can update multimodal memory when it calls the visual interface, to achieve flexible response for complex environment.
@@ 4.2 Embodied Control
Our design for embodied control capability is versatile, allowing the robot to possess a range of abilities. It is not restricted to a single type of skill but is equipped to successfully accomplish a variety of tasks in complex scenarios. Our embodied control module not only provides “tools” for the large model but also integrates multimodal environmental memory to execute the action sequence planned by the large model to complete the task.
4.3 Embodied Question Answering
Dataset. We construct a new dataset based on the simulator. First, different scenarios are constructed using the GPT-4 Turbo, with several reference examples provided. Under the guidance of prompt, the GPT-4 Turbo randomly combines the objext IDs and the location coordinates within the allowed range to produce an output. According to the output, the generation function is further called to generate corresponding items in the environment, achieving the effect of random environment construction. Next, for different types of questions, templates are designed to facilitate the large-scale generation of questions. The templates are shown in Table 1. Then, based on the randomly constructed environment, prompt is used to instruct the GPT-4 Turbo to generate questions and corresponding answers, which are combined with the scenes and question types as a piece of data in the dataset. For each scenario, three question-answer pairs are generated for each template. The dataset contains 70 generated scenes, with a total of 1050 pieces of data, all recorded in a json file. Ultimately, we conduct manual inspection and correction of the data, and ensure data balance by controlling the distribution of answers in the dataset.
type template
location What is the item on the same table as the $<obj>$?
Are the $<obj1>$ and the $<obj2>$ on the same table?
comparing Is the $<obj1>$ closer to the $<obj2>$ than to the $<obj3>$?
existence Is there any $<obj>$ in the cafe?
Is there anything in the cafe that I can use to $<do$ $something>$?
Table 1: Question Templates
Embodied Question Answering. Questions, multimodal environment memories, and newly explored information serve as inputs to the multimodal large language model GPT-4 Turbo with vision. Upon reaching a location, the robot observes the environment from four directions (front, rear, left, and right) to gather new information. Acknowledging that the plans generated by the model may not always be feasible, previously unsuccessful plans are also fed back into the model to prevent the repetition of similar invalid plans. Prompted by the designed queries, GPT-4 Turbo with vision first assesses whether the existing information is adequate for providing an accurate answer to the question. In cases of insufficiency, the model devises action plans to direct the robot in moving and interacting to acquire new observations. Then, it generates output based on the new input, thereby creating a closed loop. If the existing information is sufficient, the model synthesizes the available information to answer the question, concluding the closed loop.
@@ 4.4 Large Model
VLM. GPT-4 Turbo with vision has superior capability in both visual and language understanding compared to open-source models. In our experiments, it can form spatial understanding of complex images and offer independent judgments, thereby generating content of high quality. Therefore, we use GPT-4 Turbo with vision to implement guest seating guidance task (visual language navigation), spatial-aware dialogue interaction task (understanding the space inside the cafe and completing dialogue requirements), and control guidance task (determining which action to execute). Furthermore, we also use the model for autonomous decision-making to solve embodied question answering problems, as mentioned in Section 4.3.
@@ 5 Experimental Results
We design a plot pipeline for our solution based on reality, systematically evaluating and comparing various models for different tasks. Additionally, we also test the performance of the model on the embodied question answering dataset we construct. The details are outlined below.
@@ 5.2.4 Ablation Study
5.3 Embodied Question Answering Evaluation
In addition, we also evaluate the model on the embodied question answering task.
We randomly select 6 scenes from the dataset, with a total of 90 questions, and conduct experiments on the model in the form of single-round question answering and multi-round question answering. In the single-round question answering task, the robot starts with an empty environment memory when receiving a question and explores from scratch each time. In the multi-round question answering task, the environmental information explored to answer previous questions is available for generating answers to the current question. Each set of question-answer pairs for each scene is considered as one testing round.
The experimental results are shown in the Table 4. Metrics used to evaluate the performance of the model include average accuracy, average exploration count, average unreachable planning count, and average path length. The average accuracy(ACC) measures the proportion of correct answers among multiple questions. The average exploration count(EC) counts the number of times the robot moves to the next exploration position planned by the model, excluding the initial exploration conducted by the robot at the starting position. The average unreachable planning count(UPC) tallies the number of positions in the environment that the model plans for the robot but are inaccessible. The average path length(PL) measures the total distance in centimeters that the robot navigates in the environment from receiving the question to generating the answer.
ACC EC UPC PL
single-round 70% 22 3.5 4769.9
multi-round 72% 19.3 1 3800.7
Table 4: Comparison of results between single-round Q&A and multiple-rounds Q&A.
According to Table 4, it can be observed that in the multi-round question answering task, the model achieves better results in terms of average exploration count, average unreachable planning count, and average path length while maintaining comparable or even improved average accuracy compared to the single-round question answering task. Particularly, the average unreachable planning count is reduced by more than 70%. This result suggests the effectiveness of environment memory and historical collision information in the model’s planning process. By leveraging the environment memory and historical collision information obtained through exploration, the model can better understand the overall environment, plan action, and expedite the exploration of the required environmental information for answering questions.
5.3.1 Ablation Study
In order to further verify the effectiveness of the memory module, we present the results of the ablation experiment on the embodied question answering task. We test four models respectively: MEIA, MEIA without multimodal environment memory(MEM), MEIA without environmental language memory(ELM), and MEIA without environmental image memory(EIM). For each model, we calculate the average accuracy for different types of problems. The result is shown in Table 5.
location comparing existence total
MEIA w/o MEM 38.9% 50.0% 77.8% 56.7%
MEIA w/o ELM 41.7% 50.0% 77.8% 57.8%
MEIA w/o EIM 50.0% 55.6% 80.6% 63.3%
MEIA 55.6% 66.7% 86.1% 70.0%
Table 5: Comparison of average accuracy for different types of questions in the ablation experiment.
Obviously, MEIA performs the best among the four models. The absence of either environmental language memory or environmental image memory has an impact on the model’s performance. Furthermore, adding environmental image memory without environmental language memory does not significantly improve the model’s performance.
It can also be observed from the result that the models perform the worst on location-type questions, followed by comparing-type questions. To some extent, it reflects the limited spatial reasoning ability of MEIA, and there is room for improvement in its overall performance.
@@ 6 Conclusion
In this work, we propose the MEIA for accomplishing embodied tasks and build an embodied question answering dataset with a thousand pieces of data. In the physical environment, the agent gathers egocentric environmental information, executes embodied control and completes embodied tasks. We innovatively introduce multi-modal memory to store global memory and update it in real time, thereby integrating embodied control with large models. Experiments indicate that MEIA demonstrates the promising ability to solve various embodied tasks.